OpenStudy (anonymous):

A student says that if P(A) = P(A|B), then A and B must be independent events. Is the student correct? Explain. Give a real life example that can be represented by P(A) = P(A|B).

OpenStudy (amistre64):

what is your definition of independence?

OpenStudy (anonymous):

One event doesn't depend on the other?

OpenStudy (amistre64):

thats a little vague, how does it differ from mutually exclusive events?

OpenStudy (anonymous):

Mutually exclusive events cannot happen at the same time, but independent events can happen at the same time

OpenStudy (anonymous):

It's just that the probability of one event happening in no way affects the other happening

OpenStudy (amistre64):

correct so if the probability of an event happening, is the same for all cases ... then the probability of the event is independent of the circumstances spose there are 3 As that occur in a total of 5 Bs. P(A|B) read as, the probability of A given B, is 3/5 spose there are 6 As out of a universal set of 10 P(A|U), or simply P(A) , is 6/10 = 3/5 the probability of A is independent of the case it is a part of.

OpenStudy (amistre64):

if P(A) = P(A|B) = P(A|C) = P(A|D) = ... = P(A|K) then the probability of A is the same, for all given cases, and its value does not depend on any one specific case.

OpenStudy (anonymous):

okay, that makes sense, thank you

OpenStudy (amistre64):

good luck :)